OpenStudy (anonymous):

can someone help me get -3y= -x-3 into y=mx + b form? pleaseee!

OpenStudy (anonymous):

divide you left hand side and right hand side of the equation by -3.

OpenStudy (anonymous):

divide it??? dont you add it like something like this? -3y= -x-3 +3 +3

OpenStudy (anonymous):

is this how u divide it? -3y= -x-3 --- __ -3 -3

OpenStudy (anonymous):

-3y/-3 = -x/-3 -3/-3

OpenStudy (anonymous):

y= -x-3

OpenStudy (anonymous):

-x/-3 = x/3 -3 / -3 = 1

OpenStudy (anonymous):

OKay, so is the answer y= -x-3

OpenStudy (anonymous):

nope. the answer is y = x/3 + 1

OpenStudy (anonymous):

m=1/3, b=1

OpenStudy (anonymous):

Thank You soo much!

OpenStudy (anonymous):

Welcome :)